OpenStudy (anonymous):

the sum of the digits of a two digit number is 8.if 36 is added to the number,the digits interchange their places.find the number..

OpenStudy (anonymous):

help me guysssss

OpenStudy (helder_edwin):

Let n=a10+b. then a+b=8.

OpenStudy (helder_edwin):

n+36=b10+a

OpenStudy (anonymous):

dis is d second step???????

OpenStudy (helder_edwin):

i think u can work from here on your own.

OpenStudy (helder_edwin):

since u have two variables (the two digits) u must try to obtain two equations.

OpenStudy (helder_edwin):

well n=a10+b and n=b10+a-36

OpenStudy (helder_edwin):

so a10+b=b10+a-36 so \[ \large -9a-9b=-36 \]

OpenStudy (helder_edwin):

therefore youu have the system \[ \large \begin{cases} a+b=8\\ -9a-9b=-36 \end{cases} \]

OpenStudy (helder_edwin):

wait. there's something wrong. the system i wrote is inconsistent.

OpenStudy (anonymous):

uff

OpenStudy (anonymous):

no need i ll do it thx for ur help anyway

hartnn (hartnn):

9a-9b=-36

hartnn (hartnn):

so a10+b=b10+a-36 so −9a−9b=−36<-----* 9a-9b=-36

OpenStudy (anonymous):

hmmk

OpenStudy (helder_edwin):

u r right @hartnn

OpenStudy (helder_edwin):

so we have the system \[ \large \begin{cases} a+b=8\\ -9a+9b=36 \end{cases} \]

OpenStudy (helder_edwin):

now it is consistent!!

OpenStudy (helder_edwin):

solve this and u r done.

OpenStudy (anonymous):

okkk thx guyssssssssssssssssssss

OpenStudy (anonymous):

:)

OpenStudy (helder_edwin):

u r welcome

OpenStudy (anonymous):

thx @hartnn olsoooo thxxx

hartnn (hartnn):

welcome ^_^